### v1

# Template Matters: Understanding the Role of Instruction Templates in Multimodal Language Model Evaluation and Training

Abstract
Current multimodal language models (MLMs) evaluation and training approaches overlook the influence of instruction format, presenting an elephant-in-the-room problem. Previous research deals with this problem by manually crafting instructions, failing to yield significant insights due to limitations in diversity and scalability. In this work, we propose a programmatic instruction template generator capable of producing over 39B unique template combinations by filling randomly sampled positional synonyms into weighted sampled meta templates, enabling us to comprehensively examine the MLM’s performance across diverse instruction templates. Our experiments across eight common MLMs on five benchmark datasets reveal that MLMs have high template sensitivities with at most 29% performance gaps between different templates. We further augment the instruction tuning dataset of LLaVA-1.5 with our template generator and perform instruction tuning on LLaVA-1.5-7B and LLaVA-1.5-13B. Models tuned on our augmented dataset achieve the best overall performance when compared with the same scale MLMs tuned on at most 75 times the scale of our augmented dataset, highlighting the importance of instruction templates in MLM training. The code is available at https://github.com/shijian2001/TemplateMatters.
1 Introduction
Multimodal Language Models (MLMs) have revolutionized vision-language learning by performing visual instruction tuning on diverse, high-quality multimodal instruction data [30, 64, 17, 21, 61]. MLMs achieve unprecedented performance on various visual tasks [26, 37, 39, 58, 55]. However, previous MLM evaluation and training methods overlook a significant elephant-in-the-room problem: different instruction formats will largely influence MLMs’ performance. Although recent studies [60, 56, 32, 46] demonstrate that MLMs may produce distinct outputs when changing the instruction format (as shown in Figure 1), research on MLMs’ sensitivity to instruction formats remains largely unexplored. Previous works designed hand-crafted instructions in a limited amount, which restricts the evaluation scale, thereby weakening their conclusions, and limiting the opportunity to finetune the MLMs with augmentation on different instruction formats.
Figure 1: An example of using different instruction templates to prompt MLM without changing the original QA pairs. The instruction templates are marked in blue. Prompting MLM with different instruction templates can twist the output of MLM.
To systematically investigate the instruction sensitivity of MLMs and their impact on faithful evaluation, we propose to evaluate MLMs on Visual Question Answering (VQA) data augmented by various instruction templates without changing the meaning of the original QA pairs. To efficiently create diverse, high-quality instruction templates in sufficient quantities, we introduce a programmatic template generator that leverages diverse meta templates to produce semantically equivalent instruction templates automatically and scalably. Our approach can construct diverse instruction templates by random sampling from carefully curated word and phrase spaces to populate predefined placeholders, enabling the efficient generation of semantically consistent yet diverse instruction templates at scale. Our method can produce an extensive template space comprising 15K visual question templates and 249K choice-related templates, culminating in a comprehensive VQA instruction template space of 39B unique combinations. To effectively manage this vast template space, we use a tree-based organizational framework based on grammatical structures complemented by an efficient diversity sampling algorithm. This programmatic approach ensures the generation of instruction templates that maximize diversity across multiple dimensions, including grammatical construction, lexical choice, and symbolic representation.
We conduct a comprehensive robustness evaluation of instruction templates with our programmatic template generator, encompassing eight commonly used MLMs. Our experiment results reveal that those MLMs are highly sensitive to instruction template perturbation, with at most 29% performance gap across 100 different templates. We present the performance gap across instruction templates on the MMBench dataset in Figure LABEL:fig:banner(a). Given these results, we further introduce a simple yet effective method to improve visual instruction tuning that leverages our template generator to augment instruction datasets. We finetune two common MLMs (LLaVA-1.5-7B and LLaVA-1.5-13B) [28] using our generated diverse instruction templates and compare them with other MLMs finetuned on a larger scale (at most 75.19x than ours) of instruction-tuning datasets. Our finetuned MLMs achieve the best overall performance, demonstrating our method’s capability to improve MLMs in a data-efficient and low-cost way. We show the comparison of our 7B model to other models of similar scale on the SeedBench dataset in Figure LABEL:fig:banner(b). Our analysis further shows that compared to the original model, after finetuning with our template augmented instruction data, the model’s variance drops significantly on various out-of-domain instruction templates, which are not included in the training. Our approach not only validates the practical utility of our template generation framework but also illuminates promising directions for efficiently improving MLMs. On the other hand, our ablation studies show that models achieve the best general capabilities at a specific ratio between templates and training data, which varies with the model scale. We summarize our main contributions as follows.
  • We introduce a novel programmatic instruction template generator that enables fast and scalable generation of diverse, semantically equivalent instruction templates.
  • We evaluate the robustness of eight commonly used MLMs to instruction format variations across five benchmarks leveraging our template generator, revealing their high sensitivity to instruction format variations.
  • We propose a simple yet effective approach to enhance visual instruction tuning by augmenting the origin instruction-tuning dataset with programmatically scaling instruction templates. Our extensive experiments demonstrate its effectiveness.
2 Programmatically Scaling Instruction Templates
In this work, we propose a programmatic instruction template generator that efficiently produces a diverse array of grammatically correct and semantically consistent instruction templates without modifying the original Question-Answer (QA) pairs. Specifically, we generate instruction templates by programmatically populating placeholders in diverse meta templates with randomly sampled positional synonyms (phrases) to ensure flexibility while keeping the original meaning (Sec. 2.1). We organize our meta templates in a sentence pattern tree (Sec. 2.2), along with weighted sampling to ensure the sampling probability across all meta templates is uniformly distributed.
2.1 Meta Templates
A meta template $p_{i},i\in\{1,...,N\}$, serves as a formal blueprint for constructing instruction templates, consisting of a sequence of fixed string segments interspersed with placeholder $\langle h^{(i)}_{j}\rangle,j\in\{1,...,M_{i}\}$, where $M_{i}$ is the number of placeholders. We associate each placeholder $\langle h^{(i)}_{j}\rangle$ with a predefined set of synonyms (phrases) $s_{j}^{(i)}$. We design $s_{j}^{(i)}$ according to the semantic position of $\langle h^{(i)}_{j}\rangle$, including nouns, verbs, adjectives, or more abstract functional tokens pertinent to the context of the instruction. As illustrated in Figure 2, consider the meta template, “<verb> me <answer> to the question <related> the <image>: {question}”, where each placeholder is associated with a predefined set of positional synonyms, such as <verb> corresponds to three different candidates: “give”, “provide”, and “offer”. When generating templates, each placeholder is randomly assigned a candidate, allowing for diverse instruction templates to be produced. For example, one possible generated template is, “give me a response to the question concerning the provided image: {question}”
Figure 2: Example of the instruction template generation through a meta template.
2.2 Diverse Template Sampling
Sentence pattern tree. We build a sentence pattern tree to systematically organize our instruction template space and diversely sample our templates. We use $T=(V,E)$ to denote the sentence pattern tree, where $V$ is the set of sentence patterns and $E$ is the edge between related sentence patterns. $T$ consists of four levels, ranging from coarse-grained to fine-grained, according to the taxonomy of sentence patterns. Level 1 represents the highest level of a sentence pattern, including declarative and imperative sentences. Level 2 decomposes Level 1 into simple, complex, and compound sentences. Level 3 further breaks Level 2 into subject-predicate, subject-predicate-object, subject-subject, noun clause, gerund clause, and linking clauses. Leaves in the final level represent the meta templates belonging to the above parent nodes. We then perform weighted sampling on Level 4 according to vertice features in Level 1 to Level 3. We construct two sentence pattern trees, one consists of 24 meta templates for visual questions and another consists of 14 meta templates for choices, yielding an extensive template space encompassing 15K visual question templates and 249K choice-related templates. We present the details of our sentence pattern trees with diverse meta templates in Appendix 7.1.
Weighted sampling through sentence pattern tree. To achieve diverse sampling across the extensive template space, we implement a top-down weighted sampling approach within the sentence pattern tree. Specifically, the weight of each leaf node $\ell^{(i)}$ corresponds to the number of potential templates that can be generated by the associated meta template $p_{i}$. These weights accumulate progressively up each level of the tree, with the weight $w_{v}$ of each node $v\in V$ at any level representing the sum of weights of its descendant nodes in the next level. During sampling, we select nodes in a top-down manner, with the probability of sampling each node $v$ at a given level proportional to $w_{v}$. This process ensures that the sampling probability across all templates remains uniform, promoting diversity in generated templates while preserving the semantic consistency of each instruction template. We describe the details of our weighted sampling algorithm in Appendix 7.2.
Model BLINK MMBench SeedBench TMA MMMU
Simple Complex Simple Complex Simple Complex Simple Complex Simple Complex
Avg. 43.67 37.26 70.00 68.55 60.67 57.35 37.00 42.94 36.67 37.19
LLaVA-1.5-7B Max-Min 8.00 15.00 (+7.00) 18.00 16.00 (-2.00) 5.00 18.00 (+13.00) 14.00 26.00 (+12.00) 4.00 14.00 (+10.00)
Avg. 40.00 38.75 72.33 73.42 67.00 68.87 54.00 52.38 37.33 39.00
LLaVA-1.5-13B Max-Min 7.00 16.00 (+9.00) 3.00 12.00 (+9.00) 5.00 9.00 (+4.00) 8.00 16.00 (+8.00) 6.00 16.00 (+10.00)
Avg. 45.33 38.92 62.67 60.43 70.00 65.29 50.67 44.06 33.67 31.51
LLaVA-Next-7B Max-Min 3.00 16.00 (+13.00) 10.00 20.00 (+10.00) 2.00 18.00 (+16.00) 16.00 17.00 (+1.00) 2.00 18.00 (+16.00)
Avg. 39.67 40.72 64.67 63.47 68.33 68.76 54.67 51.53 31.00 33.23
LLaVA-Next-13B Max-Min 1.00 15.00 (+14.00) 9.00 19.00 (+10.00) 1.00 12.00 (+11.00) 5.00 21.00 (+16.00) 2.00 21.00 (+19.00)
Avg. 36.00 34.44 50.67 47.51 30.67 29.66 31.67 29.76 25.67 28.06
Qwen-VL-7B Max-Min 4.00 9.00 (+5.00) 3.00 11.00 (+8.00) 10.00 17.00 (+7.00) 9.00 19.00 (+10.00) 2.00 17.00 (+15.00)
Avg. 31.67 40.09 62.67 74.02 56.00 58.77 39.33 51.55 39.00 36.49
Qwen-VL-Chat-7B Max-Min 4.00 21.00 (+17.00) 3.00 17.00 (+14.00) 2.00 20.00 (+18.00) 8.00 17.00 (+9.00) 10.00 16.00 (+6.00)
Avg. 39.33 45.97 71.00 70.73 43.33 53.36 36.00 47.40 29.33 27.48
IDEFICS2-8B Max-Min 4.00 17.00 (+13.00) 6.00 11.00 (+5.00) 7.00 16.00 (+9.00) 8.00 20.00 (+12.00) 3.00 14.00 (+11.00)
Avg. 43.33 43.92 67.67 77.80 66.33 72.43 53.00 56.34 45.33 44.59
InternVL-Chat-1.5-24B Max-Min 6.00 24.00 (+18.00) 7.00 29.00 (+22.00) 6.00 18.00 (+12.00) 4.00 24.00 (+20.00) 1.00 17.00 (+16.00)
Table 1: Summary of our MLM evaluation results. Simple represents evaluating under three commonly used instruction templates, while Complex denotes evaluating on 100 instruction templates randomly generated from our template generator. Avg. denotes the average accuracy and Max-Min denotes the difference between best and worst accuracy across all templates. We further mark the difference of the Max-Min between Simple and Complex beside the value of Complex. The best results are marked in bold and the Max-Min values on the Complex are marked with grey. The results show that MLMs are highly sensitive to slight changes in the instruction template.
3 The Impact of Instruction Templates on MLM Performance
In this section, we leverage our programmatic instruction template generator to conduct a robust evaluation for multimodal language models (MLMs) on multiple-choice VQA tasks, which can quantitatively measure MLMs’ visual reasoning and conversational abilities.
3.1 Experiment Setup
Benchmark datasets. To comprehensively evaluate the instruction robustness of MLMs across diverse tasks and domains, we conduct our evaluation using five popular benchmark datasets: BLINK [12], SeedBench [19], MMBench [33], TaskMeAnything [60], and MMMU [59]. Each data point in the above datasets contains an image or multiple images, a question, several choices, and a correct answer. We filter these datasets to retain only the single-image samples for our evaluation. Specifically, we randomly select 100 data points for each dataset according to their category distribution, then combine each data point with (a). three simple instruction templates and (b). 100 randomly generated complex instruction templates, as shown below.
[…]
  • Complex: generated via our programmatic template generator, sampling 100 prompts from an extensive VQA template space to capture instruction format diversity.
Populating data with simple and complex templates yields two new templated datasets with 300 and 10K samples for each original dataset.
Selected models. We evaluate the performance of eight common open-source MLMs, including LLaVA-1.5-{7B, 13B} [28], LLaVA-Next-{7B, 13B} [29], Qwen-VL and Qwen-VL-Chat [3], IDEFICS2-8B [16] and InternVL-Chat-v1.5-24B [10]. We evaluate all models under the same evaluation protocol to ensure fair comparisons. Evaluating the above MLMs can give us a broad overview of open-source MLMs’ robustness to instruction formats.
Evaluation protocol. We fix the choice order according to the original dataset to eliminate this confounder and focus solely on the effects of instruction templates on model performance [63]. To retrieve answers from MLMs’ replies, we follow [60] and adopt a two-step approach. First, we apply a string-matching algorithm to determine if the model’s output matches any of three specific option representations: (1) the option identifier, e.g., (A); (2) the option content, e.g., cat; or (3) both the identifier and the name, e.g., (A) cat. If no direct match is identified, we employ a sentence-transformer [45] to calculate the embedding similarity between the model’s output and each answer option, selecting the option with the highest similarity as the predicted answer. In addition to the accuracy, we follow [46] and report the range of maximum minus minimum accuracy (Max-Min) between the highest and lowest accuracy across our generated instruction templates to quantify MLM’s sensitivity to instruction format variations.
3.2 Main Results
We evaluate eight MLMs across five datasets under two instruction template settings: Simple and Complex. For each setting, we report the average accuracy and performance range (Max-Min) across all instruction templates, as illustrated in Table 1. We present the following findings.
MLMs exhibit high sensitivity to variations in instruction templates on multiple-choice VQA tasks. As demonstrated in Table 1, most MLMs display substantial performance fluctuations under both simple and complex instruction template settings. For instance, InternVL-Chat-1.5-24B exhibits a performance difference (Max-Min) of 29% on the MMBench dataset under the complex template setting, underscoring the model’s pronounced sensitivity to instruction format variations. Furthermore, instruction format sensitivity remains consistently high regardless of the model scale. For example, a comparison between the 7B and 13B variants of both LLaVA-1.5 and LLaVA-Next reveal similarly substantial (Max-Min) values, indicating that increasing model scale doesn’t inherently reduce the sensitivity. Even after further vision instruction tuning, MLMs retain a high degree of instruction format sensitivity. Comparing Qwen-VL-7B and its instruction-tuned counterpart, Qwen-VL-Chat-7B, we observe significant (Max-Min) values across datasets for both models. This suggests that conventional vision instruction tuning can’t mitigate the instruction format sensitivity, necessitating improved vision instruction tuning.
Model comparisons may reverse depending on instruction template variations. The choices of instruction templates profoundly affect the comparative performance of MLMs, as evidenced by the variability across simple and complex instruction template settings in Table 1. For example, on the BLINK dataset, LLaVA-1.5-7B outperforms LLaVA-1.5-13B under the simple setting, while this trend reverses under the complex setting. Similarly, on the BLINK, SeedBench, and MMMU datasets, LLaVA-Next-7B achieves higher average accuracy than LLaVA-Next-13B in the simple setting, whereas LLaVA-Next-13B surpasses LLaVA-Next-7B in the complex setting. This reversal illustrates that model ranking can vary significantly based on the instruction template variations. Conclusions drawn solely from one single instruction template may lead to inaccurate comparative insights, thus underscoring the need for evaluations across diverse instruction templates to capture a comprehensive view of MLM’s performance.
Evaluations on commonly used templates tend to underestimate the performance variability of models. The results reveal a consistent pattern across models, where the performance range (Max-Min) is significantly smaller under the simple template setting than under the complex template setting. For example, in the case of the InternVL-Chat-1.5-24B model on the MMBench dataset, the (Max-Min) values for the simple and complex settings are 7 and 29, respectively, demonstrating that a limited range of instruction templates fails to capture the full extent of performance fluctuation. This disparity highlights a critical limitation in conventional MLM evaluations, as they potentially overlook the influence of instruction template diversity on model robustness. Such overlooked variability renders these evaluations less reliable for real-world applications where instruction formats are inherently diverse.
4 Visual Instruction Tuning with Diverse Instruction Templates
To tackle the issues found in Section 3, the sensitivity of MLMs to subtle changes in instruction templates, we propose a simple yet effective method that improves visual instruction tuning through a data-centric approach. Our method involves applying randomly generated instruction templates from our template generator to the original QA pairs without introducing additional data, significantly improving MLMs’ performance and reducing their sensitivity to instruction template variations. We further compare the performance of the model tuned on our method against other prominent MLMs of comparable scales (Sec. 4.2). We further conduct an ablation study to investigate how the ratio between the amount of training data and templates affects the performance of our method (Sec. 4.3).
4.1 Experiment Setup
Model # IT-Data BLINK MMB SeedB TMA MMMU Overall
S ID OOD S ID OOD S ID OOD S ID OOD S ID OOD
7B / 8B Models
Avg. 43.67 37.26 38.72 70.00 68.55 69.20 60.67 57.35 56.16 37.00 42.94 42.60 36.67 37.19 36.16 48.94
LLaVA-1.5-7B 665k Max-Min 8.00 15.00 15.00 18.00 16.00 9.00 5.00 18.00 16.00 14.00 26.00 18.00 4.00 14.00 13.00 13.93
Avg. 45.33 38.92 37.64 62.67 60.43 58.08 70.00 65.29 62.16 50.67 44.06 44.60 33.67 31.51 29.24 48.95
LLaVA-Next-7B 760k Max-Min 7.00 16.00 12.00 10.00 20.00 9.00 2.00 18.00 10.00 16.00 17.00 11.00 2.00 18.00 8.00 11.73
Avg. 36.00 34.44 34.04 50.07 47.51 47.16 30.67 29.66 28.80 31.67 29.76 30.76 25.67 28.06 28.40 34.18
Qwen-VL-7B 50M Max-Min 4.00 9.00 8.00 3.00 11.00 11.00 10.00 17.00 12.00 9.00 19.00 14.00 2.00 17.09 11.00 10.47
Avg. 31.67 40.09 40.28 62.67 74.02 75.16 56.00 58.77 58.32 39.33 51.55 51.48 39.00 36.49 36.36 50.08
Qwen-VL-Chat-7B 50M Max-Min 4.00 21.00 20.00 3.00 17.00 14.00 2.00 20.00 13.00 8.00 17.00 12.00 10.00 16.00 10.00 12.47
Avg. 39.33 45.97 46.36 71.00 70.73 70.28 43.33 53.36 54.04 36.00 47.40 46.20 29.33 27.48 28.36 47.28
IDEFICS2-8B 1.8M Max-Min 4.00 17.00 10.00 6.00 11.00 9.00 7.00 16.00 17.00 8.00 20.00 17.00 3.00 14.00 11.00 11.33
Avg. 46.33 43.19 45.44 68.67 71.66 73.20 64.33 65.13 64.16 52.00 51.78 52.64 39.33 37.46 37.32 54.18
LLaVA-1.5-7B-Base w/ Ours 665k Max-Min 5.00 13.00 2.55 10.00 12.00 8.00 3.00 11.00 6.00 4.00 22.00 10.00 9.00 11.00 6.00 8.84
13B Models
Avg. 40.00 38.75 41.20 72.33 73.42 71.24 67.00 68.87 66.92 54.00 52.38 52.24 37.33 39.00 37.20 54.13
LLaVA-1.5-13B 665k Max-Min 7.00 16.00 14.00 3.00 12.00 6.00 5.00 9.00 10.00 8.00 16.00 15.00 6.00 16.00 10.00 10.20
Avg. 39.67 40.72 38.16 64.67 63.47 63.40 68.33 68.76 66.88 54.67 51.53 47.68 31.00 33.23 33.80 51.06
LLaVA-Next-13B 760k Max-Min 1.00 15.00 13.00 9.00 19.00 15.00 1.00 12.00 11.00 5.00 21.00 14.00 2.00 21.00 10.00 11.27
Avg. 37.67 41.22 42.68 70.00 73.88 74.68 69.33 69.37 69.48 51.33 50.49 50.68 39.67 43.21 44.40 55.21
LLaVA-1.5-13B-Base w/ Ours 665k Max-Min 14.00 15.00 8.00 12.00 10.00 10.00 3.00 7.00 5.00 1.00 12.00 5.00 7.00 15.00 15.00 9.27
Table 2: Comparison of our method applied to LLaVA-1.5-7B-Base / LLaVA-1.5-13B-Base against similar-scale MLMs. Avg. denotes the average accuracy and Max-Min denotes the difference between best and worst accuracy across all templates. #IT-Data is the size of instruction tuning data the model used. S indicates the evaluation of three commonly used simple templates, ID refers to the evaluation of 100 instruction templates that our template-tuned model has encountered during training, and OOD denotes the evaluation of 25 manually crafted templates not included in our instruction template generator’s template space. The best results are marked in red bold and the second best in blue. Training with the template-augmented instruction data can boost performance across most benchmarks.
(a) 7B Models on ID templates.
(b) 7B Models on OOD templates.
(c) 13B Models on ID templates.
(d) 13B Models on OOD templates.
Figure 3: Scaling trend of the ratio of #instruction templates to #training data on the average performance across five benchmarks. There exists an optimal template-to-data ratio for MLM’s general capabilities, with stronger models requiring a smaller ratio.
BLINK
MMBench
SeedBench
TMA
MMMU
(a) Evaluation of 7B models on in-domain templates.
BLINK
MMBench
SeedBench
TMA
MMMU
(b) Evaluation of 7B models on out-of-domain templates.
BLINK
MMBench
SeedBench
TMA
MMMU
(c) Evaluation of 13B models on in-domain templates.
BLINK
MMBench
SeedBench
TMA
MMMU
(d) Evaluation of 13B models on out-of-domain templates.
Figure 5: Scaling trends of the ratio of #instruction templates to #training data on each dataset. We also show the performance spread across models and datasets. Optimal template-to-data ratios vary across datasets.
Training configurations. We trained two models based on the pretrained checkpoints: LLaVA-1.5-7B-Base and LLaVA-1.5-13B-Base, which are strong starting points for visual instruction tuning due to the open-source nature of data and models in this series. We used Low-Rank Adaptation (LoRA) [15] to train all models under the same hyperparameter settings. We used a batch size of 128 and a learning rate of $2\times 10^{-5}$ with a cosine decay schedule. The learning rate warmup ratio is set to 0.03. We used the AdamW [34] optimizer and performed fine-tuning with DeepSpeed [^1] https://github.com/microsoft/DeepSpeed at stage 3. We trained all models with 16 $\times$ A100 (40G).
Our method. We used the 665K multimodal instruction-following data [^2] https://huggingface.co/datasets/liuhaotian/LLaVA-Instruct-150K/blob/main/llava_v1_5_mix665k.json provided by the LLaVA-1.5 series. Without introducing additional data sources or training techniques, we applied instruction templates to the instruction part of the training data, resulting in a template-diversified dataset that maintains the same scale as the original. The enhanced dataset was subsequently used to finetune our pretrained LLaVA-1.5-7B-Base and LLaVA-1.5-13B-Base models.
Baseline. To establish our baseline models, we used original instruction data to perform conventional visual instruction tuning on the LLaVA-1.5-7B-Base and LLaVA-1.5-13B-Base, yielding LLaVA-1.5-7B and LLaVA-1.5-13B, which serve as our primary baselines. In addition, for the 7B model scale, we selected LLaVA-Next-7B, Qwen-VL-7B, Qwen-VL-Chat-7B, and IDEFICS2-8B as additional baselines; for the 13B model, we selected LLaVA-Next-13B as an additional baseline model. Notably, each of these additional baseline models was finetuned on a substantially larger dataset than ours.
Evaluation. We evaluated on the BLINK, MMBench, Seedbench, TaskMeAnything, and MMMU datasets. Given the computational cost associated with evaluating across multiple instruction templates, we randomly selected 100 samples from each dataset. To demonstrate the robustness of our method, we conducted evaluations under the following three instruction template settings.
(1) In-domain templates: We generated 100 templates using our template generator, which our template-tuned models have encountered during training.
(2) Out-of-domain templates: To assess the generalization ability of our method, we manually wrote 25 templates that are outside the template space of our template generator. These templates serve as a held-out set for evaluation.
(3) Commonly used simple templates: To measure the ease of use of our template-tuned model, we selected the three instruction templates from the Simple template set in Section 3.
4.2 Instruction Templates Can Improve Vision Instruction Tuning
As shown in Table 2, we compare our 7B and 13B models, trained with template-augmented instruction data, against several prominent MLMs of similar scale, revealing the following two key findings.
Template-augmented instruction data significantly enhances MLM’s performance without increasing the scale of training data. In comparison with LLaVA-1.5-7B and LLaVA-1.5-13B, which use the same pretrained models as our tuned models but rely on original instruction data, our approach of applying diverse instruction templates to the instruction part in the training data yields marked improvements across most datasets in all three evaluation settings. Furthermore, our method demonstrates superior overall performance compared to other prominent MLMs of similar scale. Remarkably, these similar-scale models were trained on much larger datasets (at most 75.19x) than ours, highlighting the effectiveness of our method in enhancing visual instruction tuning with a more efficient use of data.
Template-augmented instruction data significantly enhances MLM’s robustness to diverse instruction templates. Compared to LLaVA-1.5-7B and LLaVA-1.5-13B, our approach not only improves overall performance but also reduces the performance fluctuation range (Max-Min) across multiple instruction templates in most cases. When compared to other prominent MLMs of similar scale, our models trained with template-augmented instruction data exhibit a lower performance fluctuation range in most cases. This reduction in performance range remains stable across both in-domain (ID) and out-of-domain (OOD) instruction template settings, while counterexamples are more likely to arise with commonly used simple templates (S), given the limited use of only three evaluation templates. Notably, even when assessed using our manually written out-of-domain templates, which are outside the template space of our instruction template generator, our models frequently demonstrate a smaller performance fluctuation range. This observation underscores the effectiveness of our method in generalizing beyond the instruction templates encountered during training, rather than merely memorizing them.
4.3 Ablation on Scaling Ratio between Training Data and Templates
To investigate the impact of the ratio of instruction templates to training data (denoted as template-to-data ratio) on model performance, we created five template-augmented versions of the original 665K dataset by applying randomly sampled 100, 1K, 5K, 10K, and 15K templates. This yielded template-to-data ratios of $1.5\times 10^{-4}$, $1.5\times 10^{-3}$, $7.5\times 10^{-3}$, $1.5\times 10^{-2}$, and $2.2\times 10^{-2}$, while keeping the overall dataset size constant. Using these template-augmented datasets, we trained ten models (five with 7B parameters and five with 13B parameters) and evaluated their performance across all five benchmark datasets in both in-domain and out-of-domain template settings. Figure 3 shows the scaling curves for average performance across all datasets, while Figure 5 presents the scaling curves for each dataset. These results reveal three main findings.
MLMs perform best at specific template-to-data ratios. As shown in Figure 3, our models, which were trained with diverse instruction templates, consistently outperform models that rely on original instruction tuning data, as reflected in the average performance across five benchmark datasets. This holds across different model scales (7B and 13B), as well as for both in-domain and out-of-domain evaluation template settings, highlighting the effectiveness of our approach. Furthermore, we observe that at the 7B scale, the model achieves peak performance when the template-to-data ratio is $7.5\times 10^{-3}$, for both in-domain and out-of-domain evaluation template settings. At the 13B scale, however, the optimal ratio stabilizes at $1.5\times 10^{-4}$. The consistent scaling trends suggest the existence of a specific optimal template-to-data ratio for MLM’s general capabilities, with the model exhibiting stronger base capacity requiring a smaller optimal ratio.
Optimal template-to-data ratios vary across datasets. As shown in Figure 5, the scaling trend of the template-to-data ratio exhibits significant variability across different datasets, with the optimal ratio differing for each dataset. Furthermore, we observed that an inappropriate template-to-data ratio can lead to a decrease in performance or an increase in performance fluctuation range compared to the original model on certain datasets, revealing the limitations of our approach in specific scenarios.
Template-to-data ratio scaling trends are broadly generalizable. Whether considering the average performance in Figure 3 or the performance on individual datasets in Figure 5, both the 7B and 13B template-tuned models exhibit consistent scaling trends across in-domain and out-of-domain evaluation template settings. This consistency demonstrates that the effect of the template-to-data ratio is generalizable and doesn’t overfit the instruction templates used during finetuning.
5 Related Work
Multimodal language model.
In recent years, multimodal language models (MLMs) have advanced visual-language learning by integrating visual encoders within various pretrained large language models [48, 38, 51, 53, 4, 6, 31, 44, 7, 47, 25, 35, 20, 50, 42, 2, 49, 57]. With the increasing availability of open-sourced LLM backbones and extensive visual instruction-tuning data, models like the BLIP series [11, 22, 23, 43, 57], QwenVL series [3, 54], LLaVA series [30, 27, 29], and InternVL series [8, 9], have achieved unprecedented performance in a wide range of visual tasks [26, 37, 39, 58, 55, 1, 62]. These models, which take both visual content and language as input and output language, are now considered a new type of foundation model with exceptional visual understanding capabilities. However, these MLMs largely overlooked the significance of instruction templates of prompts, resulting in unreliable, unstable evaluation results.
Influence of template perturbation.
Recent research illustrated how prompt perturbations affect the performance and robustness of large language models (LLMs) and MLMs [14, 36, 40, 65, 13]. Liang et al. [24] performed a comprehensive examination of MLM outputs under diverse prompt designs, emphasizing the importance of systematic evaluation to ensure MLM robustness. Liu et al. [32] highlight that MLMs often produce incorrect responses when presented with nuanced, leading questions, underlining their susceptibility to prompt design variations. To solve this problem, Chatterjee et al. [5] propose a prompt sensitivity index method that captures the relative change in log-likelihood of the given prompts, making it a more reliable measure of prompt sensitivity. Some former methods [18, 41, 52] also have proposed to extend the evaluation benchmarks from a single prompt to multiple variants for each prompt. However, these former methods are all based on hand-crafted methods, which are not comprehensive enough to evaluate LLMs and MLMs. Meanwhile, most existing benchmarks, such as BLINK [12], SeedBench [19], MMBench [33], TaskMeAnything [60], and MMMU [59], still keep using a single template of the prompts for the performance evaluation.
6 Conclusion
We introduce a programmatic instruction template generator to efficiently produce diverse, high-quality instruction templates at scale, aimed at enhancing the understanding of the critical role instruction templates play in MLM evaluation and training. Using this instruction template generator, we conduct a comprehensive evaluation of MLMs’ robustness to instruction template perturbations, demonstrating the high sensitivity of MLMs to variations in instruction templates. Additionally, we propose a simple yet effective method to improve visual instruction tuning by augmenting the origin instruction-tuning dataset with programmatically scaling instruction templates, offering an efficient and cost-effective solution to improve MLMs. Our ablation studies show that models achieve the best general capabilities at a specific ratio between instruction templates and training data, which varies with the model scale.
Supplementary Material
Overview. In the supplementary material, we first detail the design and implementation of our programmatic instruction template generator in Sec. 7. Next, we present additional experiments and analysis regarding the sensitivity of MLMs to instruction templates in Sec. 8. Finally, we discuss the limitations of our work and outline directions for future work in Sec. 9.
7 Details of Instruction Template Generator
Our instruction template generator can produce an extensive template space comprising 15K visual question templates and 249K choice-related templates, culminating in a comprehensive VQA instruction template space of 39B unique combinations. Our method operates by sampling meta templates from the sentence pattern tree with a weighted sampling algorithm and then programmatically populating placeholders in meta templates with randomly sampled positional synonyms. In this section, we present the details of our sentence pattern trees with diverse meta templates (Sec. 7.1) and the weighted sampling algorithm (Sec. 7.2).
7.1 Sentence Pattern Tree with Diverse Meta Templates
(a) Sentence pattern tree with meta templates for visual questions.
(b) Sentence pattern tree with meta templates for choices.
Figure 6: Sentence pattern trees with meta templates. Each tree uses distinct colors to denote different levels. Placeholders are marked in red, while static segments are marked in black. We further mark the weight of each node (# generated templates).
We construct two sentence pattern trees, one consisting of 24 meta templates for visual questions and the other consisting of 14 meta templates for choices. To accommodate the distinct sentence structure preferences of visual questions and choice-related instruction templates, the taxonomy of these two sentence pattern trees differs slightly. We present the taxonomy and meta templates of the sentence pattern tree for visual questions in Figure 6(a) and the sentence pattern tree for choices in Figure 6(b).
7.2 Weighed Sampling Algorithm
Algorithm 1 Weight Accumulation
1:procedure AccumulateWeights($T$)
2:       for each leaf node $v$ in $T$ do
3:             $w(v)\leftarrow\text{NumTemplates}(v)$triangleright Set weight to number of potential generated templates in the leaf
4:       end for
5:       for each non-leaf node $v$ in $T$ in reverse topological order do
6:             $C\leftarrow\text{children}(v)$triangleright Retrieve children of $v$
7:             $w(v)\leftarrow\sum_{c\in C}w(c)$triangleright Sum the weights of child nodes
8:       end for
9:       return $T$triangleright Return tree with accumulated weights
10:end procedure
Algorithm 2 Weighted Sampling and Template Generation
1:procedure GenerateTemplate($T$)
2:       $v\leftarrow v_{0}$triangleright Initialize at the root node of $T$
3:       while $v$ is not a leaf node do
4:             $C\leftarrow\text{children}(v)$triangleright Retrieve child nodes of $v$
5:             $W\leftarrow\{w(c):c\in C\}$triangleright Collect weights of child nodes
6:             $v\leftarrow\text{WeightedRandomChoice}(C,W)$triangleright Select a child node based on weights
7:       end while
8:       $p\leftarrow\text{pattern}(v)$triangleright Retrieve the meta template from the selected leaf node
9:       for each placeholder $\langle h_{j}\rangle$ in $p$ do
10:             $S_{j}\leftarrow\text{synonyms}(\langle h_{j}\rangle)$triangleright Retrieve synonyms for the placeholder
11:             $s_{j}\leftarrow\text{UniformRandomChoice}(S_{j})$triangleright Randomly select a synonym
12:             Replace $\langle h_{j}\rangle$ in $p$ with $s_{j}$triangleright Substitute placeholder with synonym
13:       end for
14:       return $p$triangleright Return the constructed instruction template
15:end procedure
To ensure diverse meta template sampling, we propose a weighted sampling algorithm within the sentence pattern tree that guarantees a uniform probability distribution across all meta templates.
We begin by implementing an automatic weight accumulation algorithm for the sentence pattern tree. Each leaf node (meta template) is assigned a weight corresponding to the number of templates it can potentially generate. These weights are then propagated upward, with the weight of each non-leaf node calculated as the sum of the weights of its child nodes. The detailed procedure for this algorithm is outlined in Algorithm 1.
After constructing the weighted sentence pattern tree, we perform top-down sampling, selecting nodes at each layer with probabilities proportional to their weights. Upon reaching a leaf node, we programmatically populate the placeholders in the corresponding meta template with randomly selected positional synonyms, resulting in grammatically correct and diverse instruction templates. We present the details of the procedure in Algorithm 2.
8 Additional Experiments on MLM’s Sensitivity to Instruction Templates
Figure 7: Heat map illustrating the performance variations of eight MLMs on the BLINK dataset across ten instruction templates (selected from the Complex templates set in Sec. 3). The darker the color, the better the performance. No single instruction template performs optimally for all MLMs.
In this section, we explore whether a universally effective instruction template for most MLMs exists. To this end, we analyze the performance of eight prominent MLMs on the BLINK dataset using ten instruction templates selected from the Complex templates set as described in Section 3. A heat map is presented in Figure 7 to illustrate the performance variations, where darker shades correspond to superior performance.
The results reveal substantial performance variability across MLMs for the same instruction template, as indicated by the diverse color gradients within each column of the heat map. This variability highlights a critical observation: no single instruction template consistently performs optimally for all MLMs. This lack of universality implies that each model exhibits distinct sensitivities and preferences toward different instruction template, which complicates the task of designing or selecting a universally effective template.
The observed variations have important implications for the instruction template design and evaluation of MLMs. Specifically, they underscore the limitations of a one-size-fits-all approach to instruction optimization. Efforts to identify an ideal instruction template through a single, static search are unlikely to yield universally effective results. Instead, tailored strategies that consider the specific characteristics and requirements of individual MLMs can be necessary to achieve optimal performance.
9 Discussion
9.1 Limitation
Designing the template space requires manual effort.
The development of meta templates and the association of placeholders with semantically equivalent synonyms demand significant manual intervention. Despite the automation of template generation, ensuring semantic consistency and grammatical correctness across diverse templates is labor-intensive.
Evaluation across multiple instruction templates is cost-prohibitive.
Evaluating MLMs with extensive template spaces incurs high computational costs due to the increased number of evaluations per dataset. This limits the scalability of testing, especially for large datasets or when comparing multiple models. The high costs associated with such exhaustive evaluations often necessitate trade-offs, limiting the breadth of experimentation and potentially overlooking optimal template configurations.
An imbalance in the template-to-data ratio during training can degrade model performance on specific datasets.
The results in Sec. 4.3 indicate that models achieve peak performance at specific template-to-data ratios, which vary based on model scale and dataset. Disproportionate scaling of either templates or data can lead to performance variability and generalization challenges.
9.2 Future Work
Budget-constrained instruction template optimization tailored to specific models and tasks.
The findings in Sec. 8 indicate that no universal optimal instruction template exists across all models. However, for a specific model and dataset, it is practical and valuable to identify the most effective instruction template from a large pool of predefined options within a constrained computational budget. Our future work will explore developing efficient methods for optimizing instruction templates to enhance task-specific model performance.
Enhancing the generalization of template-augmented training.
The conclusions present in Sec. 4.3 highlight the limitations of our approach when faced with an imbalanced template-to-data ratio. To address this, our future research will explore developing advanced techniques to enhance the generalization capabilities of our template augmentation methods, ensuring its robustness across diverse scenarios and datasets.